OpenStudy (anonymous):

The volume of a rectangular prism is represented by the function x^3 + 11x^2 + 20x − 32. The width of the box is x − 1 while the height is x + 8. Find the expression representing the length of the box. A. x − 6 B. x − 4 C. x + 4 D. x + 6

OpenStudy (da_scienceman):

what is the formula for the volume of a rectangular prism? The info from your question implies the volume = lwh (w=width, h=height, l=length) equals your expression. So I will multiply the expression for the height and width then divide by the expression for the volume. The answer gives me the length of the prism!

OpenStudy (jdoe0001):

\(\bf \textit{volume of a cube}=length\cdot width\cdot height\\ \quad \\ length\cdot width\cdot height=length\cdot width\cdot height\\ \quad \\\implies \cfrac{length\cdot width\cdot height}{length\cdot width}=height\)

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

rather..... \(\bf \bf \textit{volume of a cube}=length\cdot width\cdot height\\ \quad \\ length\cdot width\cdot height=length\cdot width\cdot height\\ \quad \\\implies \cfrac{length\cdot width\cdot height}{height\cdot width}=length\)

OpenStudy (jdoe0001):

\(\bf \textit{volume of a cube}= x^3 + 11x^2 + 20x − 32\\ \quad \\ width=(x-1)\qquad \qquad height=(x+8)\qquad length = l\\ \quad \\ \quad \\ x^3 + 11x^2 + 20x − 32=(x-1)(x+8)l\\ \quad \\\implies \cfrac{x^3 + 11x^2 + 20x − 32}{(x-1)(x+8)}=l\)

OpenStudy (da_scienceman):

No I think you will find the expression by dividing \[x^{3} + 11x ^{2} + 20x - 32 \] by \[x ^{2}+7x -8\]

OpenStudy (da_scienceman):

the answer will give u the required expression!